An object wtih mass m has moment of inertia I about its center of mass. If we take the moment of inertia at a distance 3d from the center of mass, how does I change?


1
Expert's answer
2022-01-10T16:56:09-0500

"I'=I+m(3d)^2=I+9md^2."